2.0 mol of PCl5 were introduced in a vessel of 5.0 L capacity at a particular temperature . At equilibrium PCl5 was found to be 35% dissociated into PCl3 and Cl2 . The value of Kc for the reaction is

2.0 mol  of PCl5  were introduced in a vessel  of 5.0 L capacity  at a particular temperature . At   equilibrium PCl5 was found to be 35% dissociated into PCl3 and Cl2 . The value of Kc for the reaction is

Grade:9

1 Answers

Vikas TU
7 years ago
                     PCl5     ---------->         PCl3         +        Cl2
t=0                 0.4                             0                        0
t=teq.           0.4-x                            x                         x
At eqm., (0.4 -x) = 0.65*0.4
Or
x = 0.4 – 0.4*0.65
   =0.4(1-0.65)
   =0.4*0.35
   = 0.14
 
Kc = 0.14*0.14/0.26  
     = 0.0753846